# General Coded Computing in a Probabilistic Straggler Regime

### v1

@@ I Introduction
Notations: In this paper, bold uppercase letters (e.g., $\mathbf{A}$) are used to denote matrices, while bold lowercase letters (e.g., $\mathbf{x}$) represent vectors. Coded matrices and vectors are indicated with a tilde, such as $\tilde{\mathbf{x}}$ and $\tilde{\mathbf{A}}$. The set $\{1,2,\ldots,n\}$ is denoted by $[n]$, and the size of a set $S$ is written as $|S|$. The $i$-th component of a vector-valued function $\mathbf{f}$ is expressed as $f_{i}(\cdot)$. The first, second, and $k$-th order derivatives of a scalar function $f$ are represented as $f^{\prime}$, $f^{\prime\prime}$, and $f^{(k)}$, respectively. The $\ell_{2}$-norm of a vector $\mathbf{x}$ is denoted by $\left\lVert\mathbf{x}\right\rVert_{2}$, while the $L_{p}$-norm of a function $f(\cdot)$ over the interval $\Omega$ is written as $\left\lVert f\right\rVert_{{L}^{p}\left(\Omega\right)}$. The space $\mathcal{H}^{2}\left(\Omega;\mathbb{R}^{m}\right)$ corresponds to the reproducing kernel Hilbert space (RKHS) of second-order Sobolev functions for vector-valued functions of dimension $m$ over $\Omega$. Additionally, $\mathcal{C}^{2}\left(\Omega;\mathbb{R}^{m}\right)$ represents the space of all vector-valued functions of dimension $m$ that with a second derivative over the interval $\Omega$. For one dimension, we simply use $\mathcal{H}^{2}\left(\Omega\right)$ and $\mathcal{C}^{2}\left(\Omega\right)$.
@@ II-B Coded Computing Frameworks for General Computing Function
[…]
@@ III Problem Formulation
$\mathcal{L}(\mathbf{\hat{f}})$ $:=\mathop{\mathrm{E}}_{\mathcal{F}\sim F_{N}}\left[\frac{1}{K}\sum^{K}_{k=1}\left\lVert\mathbf{\hat{f}}(\mathbf{x}_{k})-\mathbf{f}(\mathbf{x}_{k})\right\rVert^{2}_{2}\right]=\mathop{\mathrm{E}}_{\mathcal{F}\sim F_{N}}\left[\frac{1}{K}\sum^{K}_{k=1}\left\lVert\mathbf{u}_{\textrm{dec}}(\alpha_{k})-\mathbf{f}(\mathbf{x}_{k})\right\rVert^{2}_{2}\right].$ (5)
Our goal is to derive an upper bound for $\mathcal{L}(\mathbf{\hat{f}})$ and compare the performance of BACC and LeTCC schemes in probabilistic configuration of stragglers.
@@ IV Main Results
Consider the LeTCC framework consisting of $N$ servers, where $u_{\textrm{enc}}(\cdot),u_{\textrm{dec}}(\cdot)\in\mathcal{H}^{2}\left(\Omega\right)$ and $\lambda_{\textrm{dec}}\leqslant N^{-4}$. Let $f(\cdot)$ be a $\nu$-Lipschitz function satisfying $|f^{\prime\prime}|\leq\eta$ for some $\eta>0$. Assume that each server becomes a straggler with probability $p:=1-q>0$. If $\frac{\Delta_{\textrm{max}}}{\Delta_{\textrm{min}}}\leqslant B$ for some constant $B>0$, then for any $\delta\in(0,1)$, there exist constants $C_{l}$ and $n_{0}$ such that for $N>n_{0}$, with probability at least $1-\delta$, the following holds:
@@
Consider the BACC framework consisting of $N$ servers with $f\in\mathcal{C}^{2}\left(\Omega\right)$. Assume that each server is a straggler with probability of $p:=1-q>0$. If there is a constant $B$ such that $\frac{\Delta_{\textrm{max}}}{\Delta_{\textrm{min}}}\leqslant B$, then for any $\delta\in(0,1)$, there exist constants $C_{b}$ and $n_{0}$ such that for $N>n_{0}$ with a probability of at least $1-\delta$, we have:
@@ V Proof Sketch
In this section, we provide a proof sketch for the main theorems, with the formal proofs available in [15]. We first bound the average approximation error for each of LeTCC and BACC schemes.
@@
Since $f(\cdot)$ is $\nu$-Lipschitz, we can derive an upper bound for $\mathcal{L}_{\textrm{enc}}(\hat{f})$:
@@
Next, we drive an upper bound for $\mathcal{L}_{\textrm{dec}}(\hat{f})$. Defining $h(z):=u_{\textrm{dec}}(z)-f(u_{\textrm{enc}}(z))$, we obtain $\mathcal{L}_{\textrm{dec}}(\hat{f})\leqslant 2\mathrm{E}_{\mathcal{F}\sim F_{N}}\left[\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}\right].$ Given that $f(\cdot)$ has bounded first and second derivatives, we can prove that $h\in\mathcal{H}^{2}\left(\Omega\right)$. Applying Sobolev interpolation inequalities [16], we can derive an upper bound for $\left\lVert h\right\rVert^{2}_{{L}^{\infty}\left(\Omega\right)}$:
@@
where $\mu:=\max_{1\leq i\leq|\mathcal{F}|-2}\min\left\{\frac{\beta_{j_{i+1}}-\beta_{j_{i}}}{\beta_{j_{i}}-\beta_{j_{i-1}}},\frac{\beta_{j_{i+1}}-\beta_{j_{i}}}{\beta_{j_{i+2}}-\beta_{j_{i+1}}}\right\}$ and :=$\mathcal{F}:=\{\beta_{j_{1}},\dots,\beta_{j_{|\mathcal{F}|}}\}$. Applying Max-Min inequality, we have: $\mu\leqslant\min\left\{\max_{1\leq i\leq|\mathcal{F}|-2}\frac{\beta_{j_{i+1}}-\beta_{j_{i}}}{\beta_{j_{i}}-\beta_{j_{i-1}}},\max_{1\leq i\leq|\mathcal{F}|-2}\frac{\beta_{j_{i+1}}-\beta_{j_{i}}}{\beta_{j_{i+2}}-\beta_{j_{i+1}}}\right\}\leqslant\frac{\Delta^{\mathcal{F}}_{\textrm{max}}}{\Delta^{\mathcal{F}}_{\textrm{min}}}.$ Thus:
@@
([26, Theorem 2]) Let $R_{\mathcal{F},N}$ be the maximal sub-sequence length of consecutive ones or length of longest run in an i.i.d. Bernoulli sequence $B_{1},\dots,B_{N}$ with $\mathrm{E}[B_{i}]=p=1-q$ for $i\in[N]$. Let $\theta=\frac{\pi^{2}}{\ln\left(\frac{1}{p}\right)}$. Then:

### v2

@@ I Introduction
Notations: In this paper, bold uppercase letters (e.g., $\mathbf{A}$) are used to denote matrices, while bold lowercase letters (e.g., $\mathbf{x}$) represent vectors. Coded matrices and vectors are indicated with a tilde, such as $\tilde{\mathbf{x}}$ and $\tilde{\mathbf{A}}$. The set $\{1,2,\ldots,n\}$ is denoted by $[n]$, and the size of a set $S$ is written as $|S|$. The $i$-th component of a vector-valued function $\mathbf{f}$ is expressed as $f_{i}(\cdot)$. The first, second, and $k$-th order derivatives of a scalar function $f$ are represented as $^{\prime}, f′′f^{\prime\prime}$, and $f^{(k)}$, respectively. The $\ell_{2}$-norm of a vector $\mathbf{x}$ is denoted by $\left\lVert\mathbf{x}\right\rVert_{2}$, while the $L_{p}$-norm of a function $f(\cdot)$ over the interval $\Omega$ is written as $\left\lVert f\right\rVert_{{L}^{p}\left(\Omega\right)}$. The space $\mathcal{H}^{2}\left(\Omega;\mathbb{R}^{m}\right)$ corresponds to the reproducing kernel Hilbert space (RKHS) of second-order Sobolev functions for vector-valued functions of dimension $m$ over $\Omega$. Additionally, $\mathcal{C}^{2}\left(\Omega;\mathbb{R}^{m}\right)$ represents the space of all vector-valued functions of dimension $m$ that with a second derivative over the interval $\Omega$. For one dimension, we simply use $\mathcal{H}^{2}\left(\Omega\right)$ and $\mathcal{C}^{2}\left(\Omega\right)$.
@@ II-B Coded Computing Frameworks for General Computing Function
$\mathbf{u}_{\textrm{enc}}(\cdot)$ $=\underset{\mathbf{u}\in\mathcal{H}^{2}\left(\Omega;\mathbb{R}^{d}\right)}{\operatorname{argmin}}\frac{1}{|\mathcal{F}|}\sum^{K}_{k=1}\left\lVert\mathbf{u}\left(\alpha_{k}\right)-\mathbf{x}_{k}\right\rVert^{2}_{2}+\lambda_{\textrm{enc}}\sum^{d}_{j=1}\int_{\Omega}\left(u_{j}^{\prime\prime}(t)\right)^{2}\,dt,$ (3)
$\mathbf{u}_{\textrm{dec}}(\cdot)$ $\underset{\mathbf{u}\in\mathcal{H}^{2}\left(\Omega;\mathbb{R}^{m}\right)}{\operatorname{argmin}}\frac{1}{|\mathcal{F}|}\sum_{v\in\mathcal{F}}\left\lVert\mathbf{u}\left(\beta_{v}\right)-\mathbf{f}\left(\mathbf{u}_{\textrm{enc}}\left(\beta_{v}\right)\right)\right\rVert^{2}_{2}+\lambda_{\textrm{dec}}\sum^{m}_{j=1}\int_{\Omega}\left(u_{j}^{\prime\prime}(t)\right)^{2}\,dt,$ (4)
@@ III Problem Formulation
[…]
@@ IV Main Results
Consider the LeTCC framework consisting of $N$ servers, where $u_{\textrm{enc}}(\cdot),u_{\textrm{dec}}(\cdot)\in\mathcal{H}^{2}\left(\Omega\right)$ and $\lambda_{\textrm{dec}}\leqslant N^{-4}$. Let $f(\cdot)$ be a $\nu$-Lipschitz function satisfying $|f^{\prime\prime}|\leq\eta$ for some $\eta>0$. Assume that each server becomes a straggler with probability p:=1q>0p:=1-q>0. If $\frac{\Delta_{\textrm{max}}}{\Delta_{\textrm{min}}}\leqslant B$ for some constant $B>0$, then for any $\delta\in(0,1)$, there exist constants $C_{l}$ and $n_{0}$ such that for $N>n_{0}$, with probability at least $1-\delta$, the following holds:
@@
Consider the BACC framework consisting of $N$ servers with $f\in\mathcal{C}^{2}\left(\Omega\right)$. Assume that each server is a straggler with probability of p:=1q>0p:=1-q>0. If there is a constant $B$ such that $\frac{\Delta_{\textrm{max}}}{\Delta_{\textrm{min}}}\leqslant B$, then for any $\delta\in(0,1)$, there exist constants $C_{b}$ and $n_{0}$ such that for $N>n_{0}$ with a probability of at least $1-\delta$, we have:
@@ V Proof Sketch
In this section, we provide a proof sketch for the main theorems, with the formal proofs available in Appendix A. We first bound the average approximation error for each of LeTCC and BACC schemes.
@@
[…]
@@
[…]
@@
where $\mu:=\max_{1\leq i\leq|\mathcal{F}|-2}\min\left\{\frac{\beta_{j_{i+1}}-\beta_{j_{i}}}{\beta_{j_{i}}-\beta_{j_{i-1}}},\frac{\beta_{j_{i+1}}-\beta_{j_{i}}}{\beta_{j_{i+2}}-\beta_{j_{i+1}}}\right\}$ and $\mathcal{F}:=\{\beta_{j_{1}},\dots,\beta_{j_{|\mathcal{F}|}}\}$. Applying Max-Min inequality, we have: $\mu\leqslant\min\left\{\max_{1\leq i\leq|\mathcal{F}|-2}\frac{\beta_{j_{i+1}}-\beta_{j_{i}}}{\beta_{j_{i}}-\beta_{j_{i-1}}},\max_{1\leq i\leq|\mathcal{F}|-2}\frac{\beta_{j_{i+1}}-\beta_{j_{i}}}{\beta_{j_{i+2}}-\beta_{j_{i+1}}}\right\}\leqslant\frac{\Delta^{\mathcal{F}}_{\textrm{max}}}{\Delta^{\mathcal{F}}_{\textrm{min}}}.$ Thus:
@@
([25, Theorem 2]) Let $R_{\mathcal{F},N}$ be the maximal sub-sequence length of consecutive ones or length of longest run in an i.i.d. Bernoulli sequence $B_{1},\dots,B_{N}$ with $\mathrm{E}[B_{i}]=p=1-q$ for $i\in[N]$. Let $\theta=\frac{\pi^{2}}{\ln\left(\frac{1}{p}\right)}$. Then: